OpenStudy (anonymous):

what is the length of the transeverse axis (x-2)^2/36-(y+1)^2/64=1 HELP PLEASE

OpenStudy (anonymous):

20

OpenStudy (anonymous):

x axis 2a = 2 * 6 = 12 y axis 2b = 2 * 8 = 16

OpenStudy (anonymous):

the equation for the directrix of this parabola is x=-1. x=1/4y^2 true or false

OpenStudy (anonymous):

Just post your questions, I'm going to reply soon, right now my PC ask me to restart :(